# Extensions of $\mathcal{KL}$ and Lyapunov Functions for Discrete-time Dynamical System Peaks Analysis

### v3

@@ 4.2 Revisiting $\mathcal{KL}$ functions : the class $\mathcal{KL}_{\rm gen}$
Suppose that there exists a $\mathcal{KL}_{\rm gen}$-$\operatorname{Fin}\left(X^{\mathrm{in}}\right)$ upper bound $(\gamma,\theta)$ for $(X^{\mathrm{in}},T,\varphi)$. As $s\mapsto\gamma(s,t)$ is increasing for all $t\geq 0$, we have for all $x\in X^{\mathrm{in}}$, for all $k\in\mathbb{N}$, $\gamma(\theta(x),k)\leq\gamma(\overline{\theta_{X^{\mathrm{in}}}},k)$ and thus $\nu_{k}=\sup\{\varphi(T^{k}(x)):x\in X^{\mathrm{in}}\}\leq\gamma(\overline{\theta_{X^{\mathrm{in}}}},k)$ for all $k\in\mathbb{N}$. Now, taking the limsup on both sides leads to $\limsup_{k\mapsto+\infty}\nu_{k}\leq\inf_{k\in\mathbb{N}}\gamma(\overline{\theta_{X^{\mathrm{in}}}},k)\leq\gamma(\overline{\theta_{X^{\mathrm{in}}}},0)<+\infty$ as $t\mapsto\gamma(s,t)$ is decreasing for all $s\in\mathbb{R}$.
Suppose that $\nu=(\nu_{k})_{k\in\mathbb{N}}\in\Lambda$. From Prop. 8, $\Gamma(\nu)$ is nonempty and let $(h,\beta)$ be in $\Gamma(\nu)$. Then, we define for all $s\in\mathbb{R}$ and $t\in\mathbb{R}_{+}$, $\gamma(s,t):=\min\{s,h(\beta^{t})\}$ and for all $x\in\mathbb{R}^{d}$, $\theta(x):=\sup_{k\in\mathbb{N}}\varphi(T^{k}(x))$. It is clear that $\gamma\in\mathcal{KL}_{\rm gen}$ as $\beta\in(0,1)$. Now as $\nu\in\Lambda$, $\nu_{\rm opt}=\sup_{n\in\mathbb{N}}\nu_{n}\in\mathbb{R}$. Moreover, we have $\nu_{\rm opt}=\sup_{n\in\mathbb{N}}\sup_{x\in X^{\mathrm{in}}}\varphi(T^{n}(x))=\sup_{x\in X^{\mathrm{in}}}\sup_{n\in\mathbb{N}}\varphi(T^{n}(x))=\overline{\theta_{X^{\mathrm{in}}}}$. By definition of $\theta$, we have for all $x\in X^{\mathrm{in}}$, for all $k\in\mathbb{N}$, $\varphi(T^{k}(x))\leq\theta(x)$. By definition of $(h,\beta)$, we have for all $k\in\mathbb{N}$, $\nu_{k}\leq h(\beta^{k})$. This is the same as for all $x\in X^{\mathrm{in}}$ and for all $k\in X^{\mathrm{in}}$, $\varphi(T^{k}(x))\leq h(\beta^{k})$. We conclude that $\varphi(T^{k}(x))\leq\min\{\theta(x),h(\beta^{k})\}$ for all $x\in X^{\mathrm{in}}$ and for all $k\in\mathbb{N}$. Note that for a function $\gamma$ constructed as below, for all $s\in\mathbb{R}$, $\inf_{t\geq 0}\gamma(s,t)=\inf_{t\geq 0}\min\{s,h(\beta^{t})\}=\min\{s,h(0)\}$. This follows from the continuity of $h$, the fact $h$ is strictly increasing and $\beta\in(0,1)$.
@@
If moreover, ${\mathrm{R}}_{\nu}^{s}\neq\emptyset$, then from Lemma 10, there exists $h\in\mathbb{\Gamma}_{\rm fun}(\nu)$ useful for $\nu$. We could define $\gamma$ using this function $h$. Now, let $k\in\mathcal{S}(\nu,h)$. As $\nu_{k}\leq\nu_{\rm opt}=\overline{\theta_{X^{\mathrm{in}}}}$, we conclude that $\inf_{t\geq 0}\gamma(\overline{\theta_{X^{\mathrm{in}}}},t)=\min\{\overline{\theta_{X^{\mathrm{in}}}},h(0)\}=h(0)$ which ends the proof. ∎
Example 4.
Let us come back to the analysis of the peaks computation problem of Subsection 2.2. We propose to compute a useful $\mathcal{KL}_{\rm gen}$-$\operatorname{Fin}\left(X^{\mathrm{in}}\right)$ upper bound $(\gamma,\theta)$ for the triplet $(X^{\mathrm{in}},T,\varphi)$ defined at Subsection 2.2.
Following the proof of Th 5, one can define the $\mathcal{KL}_{\rm gen}$ function as $(s,t)\mapsto\min\{s,h(\beta^{t})\}$ where $(h,\beta)$ is a pair proposed at Example 2 and $\theta\in\operatorname{Fin}\left(X^{\mathrm{in}}\right)$ as $\mathbb{R}^{2}\ni x\mapsto\sup_{n\in\mathbb{N}}\varphi(T^{n}(x))$.
Another possible pair consists in the $\mathcal{KL}_{\rm gen}$ function: ×$\mathbb{R}\times\mathbb{R}_{+}\ni(s,t)\mapsto\min\{s,(p^{2}(n_{0}+1)(3(t+1))^{-1}\}$ and the constant function in $\operatorname{Fin}\left(X^{\mathrm{in}}\right)$: $\mathbb{R}^{2}\ni x\mapsto p^{2}/3$. We recall that $n_{0}$ is the smallest integer $n$ such that $\varphi_{n}(\alpha)<0$. We can easily check that $\varphi(T^{k}(x))\leq\min\{p^{2}/3,(p^{2}(n_{0}+1)(3(k+1))^{-1}\}$ for all $x\in X^{\mathrm{in}}$ and $k\in\mathbb{N}$. Moreover, as $\inf_{t\geq 0}\min\{p^{2}/3,(p^{2}(n_{0}+1)(3(t+1))^{-1}\}=0$, there exists $k\in\mathbb{N}$ such that $\nu_{k}>\inf_{t\geq 0}\min\{p^{2}/3,(p^{2}(n_{0}+1)(3(t+1))^{-1}\}$.
@@
In the proof of Th. 5, we exploit Prop. 8 and Prop. 10 which link the nonemptiness of the sets $\mathbb{\Gamma}(u)$, $\mathcal{S}(u,h)$ and ${\mathrm{R}}_{u}^{s}$. We can reformulate Th 5 as Corollary 3.
@@
In Th 5 and in Corollary 3, we construct a useful $\mathcal{KL}_{\rm gen}$-$\operatorname{Fin}\left(X^{\mathrm{in}}\right)$ upper bound for $(X^{\mathrm{in}},T,\varphi)$ from a useful $(h,\beta)\in\mathbb{\Gamma}(\nu)$. Now, the question is to know whether is it possible to construct a useful $(h,\beta)\in\mathbb{\Gamma}(\nu)$ from a useful $\mathcal{KL}_{\rm gen}$-$\operatorname{Fin}\left(X^{\mathrm{in}}\right)$ upper bound for $(X^{\mathrm{in}},T,\varphi)$. This is exactly what was done at Th 4 from the Sontag’s result. Thus, we provide a positive answer. The result is based on an extension of the Sontag’s result for the class $\mathcal{KL}_{\rm gen}$. This permits to fix the value $\beta\in(0,1)$ to only consider problem on functions in $\Omega([0,1])$.
@@
Let $f:\mathbb{R}_{+}\to\mathbb{R}$ be a decreasing function and let $m\geq\inf_{\mathbb{R}_{+}}f$ ($m$ can be equal to $\infty$- ∞ if $\inf_{\mathbb{R}_{+}}f$ is). There exists a function $g:\mathbb{R}_{+}\mapsto\mathbb{R}$ strictly decreasing and continuous such that:
@@
Now assume that $f(0)>m\geq\inf_{\mathbb{R}_{+}}f$. We denote for all couple of pairs $\{(x,y),(z,t)\}$ where $x,y,z,t\in\mathbb{R}$ by $\mathrm{Aff}\{(x,y),(z,t)\}$ the unique affine function $\ell:\mathbb{R}\mapsto\mathbb{R}$ such that $\ell(x)=y$ and $\ell(z)=t$. We define the functions $h_{0}$ and $h_{n}$ for all $n\in\mathbb{N}^{*}$ as follows :
@@
$F(x)=\sum_{n\in\mathbb{N}}F_{n}(x)+\dfrac{1}{x+1}\text{ and note that}\sum_{n\in\mathbb{N}}F_{n}(x)=\left\{\begin{array}[]{lr}f(0)&\text{ if }x\in[0,1]\\ f(\lfloor x\rfloor-1)&\text{ if }x\in[\lfloor x\rfloor,\lfloor x\rfloor+1/2]\\ \ell\left(x\right)&\text{ if }x\in(\lfloor x\rfloor+1/2,\lfloor x\rfloor+1)\end{array}\right.$
@@
The function $\sum_{n\in\mathbb{N}}F_{n}$ is clearly continuous and decreasing (as $f$ does) on $\mathbb{R}_{+}$. The function $\sum_{n\in\mathbb{N}}F_{n}$ is not strictly decreasing but adding $x\mapsto(x+1)^{-1}$ makes $F$ to be strictly decreasing and continuous. We also have $f\leq\sum_{n\in\mathbb{N}}F_{n}$ as $f$ is decreasing. As $x\mapsto(x+1)^{-1}$ is strictly positive we have $f\leq F$. Finally, we have $\inf_{\mathbb{R}_{+}}F=\inf_{\mathbb{R}_{+}}f$. Indeed, for all $n\in\mathbb{N}^{*}$ and all $x\in[n,n+1)$, $F_{n}\leq f(n-1)$, it implies that $\inf_{\mathbb{R}_{+}}F=\inf_{x\in\mathbb{R}_{+}}\inf_{n\in\mathbb{N}^{*}}F_{n}(x)\leq\inf_{n\in\mathbb{N}^{*}}f(n-1)=\inf_{\mathbb{R}_{+}}f$ (as $f$ decreases). The reverse inequality holds as $f\leq F$.
The function $F$ does not necessarily satisfy $\inf_{\mathbb{R}_{+}}h=m$. Hence, we define for all $x\in\mathbb{R}_{+}$:
@@
Suppose that $m>\inf_{\mathbb{R}_{+}}f$. The function $g$ is clearly strictly decreasing and continuous and satisfies $f\leq g$ as $F$ does. As $m>\inf_{\mathbb{R}_{+}}f=\inf_{\mathbb{R}_{+}}F$, there exists $x_{0}$ such that $F(x)<m$ for all $x\geq x_{0}$ and then $g(x)=(x+1)^{-1}+m$ for all $x\geq x_{0}$. This implies that $\inf_{\mathbb{R}_{+}}g=m$. ∎
@@
By definition, $t\mapsto\gamma(s,t)$ is decreasing and applying Lemma 2, there exists a function $\sigma:\mathbb{R}_{+}\mapsto\mathbb{R}$ strictly decreasing and continuous such that $\gamma(s,t)\leq\sigma(t)$ for all $t\geq 0$ and $\inf_{\mathbb{R}_{+}}\sigma=m$. Let us define for all $r\in(0,1]$, $h(r)=\sigma(-\ln r)$ and $\lim_{x\to 0^{+}}h(x)=\lim_{y\to+\infty}\sigma(y)=m$. Then defining $h(0)=m$ makes $h$ continuous at 0. Finally, $h$ belongs to $\Omega([0,1])$. ∎
Proof of Th. 6.
@@
By assumption, there exist a $\mathcal{KL}_{\rm gen}$-$\operatorname{Fin}\left(X^{\mathrm{in}}\right)$ upper bound $(\gamma,\theta)$ for $(X^{\mathrm{in}},T,\varphi)$. Let $x\in X^{\mathrm{in}}$ and $k\in\mathbb{N}$. This implies that $\varphi(T^{k}(x))\leq\gamma(\overline{\theta_{X^{\mathrm{in}}}},k)$. By Lemma 3, there exists $h\in\Omega([0,1])$ such that $\gamma(\overline{\theta_{X^{\mathrm{in}}}},k)\leq h(e^{-k})$ for all $k\in\mathbb{N}$. This leads to $(h,e^{-1})\in\mathbb{\Gamma}(\nu)$. If there exists $k\in\mathbb{N}$ such that $\inf_{t\in\mathbb{R}}\gamma(\overline{\theta_{X^{\mathrm{in}}}},t)<\nu_{k}$, we can choose, as it is stated in Lemma 3, $h$ such that $h(0)=m\in(\inf_{t\in\mathbb{R}}\gamma(\overline{\psi},t),\nu_{k})$. ∎

### v4

@@ 4.1 Revisiting $\mathcal{KL}$ functions : the class $\mathcal{KL}_{\rm gen}$
Suppose that there exists a $\mathcal{KL}_{\rm gen}$-$\operatorname{Fin}\left(X^{\mathrm{in}}\right)$ upper bound $(\gamma,\theta)$ for $(X^{\mathrm{in}},T,\varphi)$. As $s\mapsto\gamma(s,t)$ is increasing for all $t\geq 0$, we have for all $x\in X^{\mathrm{in}}$, for all $k\in\mathbb{N}$, $\gamma(\theta(x),k)\leq\gamma(\overline{\theta_{X^{\mathrm{in}}}},k)$ and thus $\nu_{k}=\sup\{\varphi(T^{k}(x)):x\in X^{\mathrm{in}}\}\leq\gamma(\overline{\theta_{X^{\mathrm{in}}}},k)$ for all $k\in\mathbb{N}$. Now, taking the limsup on both sides leads to
$\limsup_{k\mapsto+\infty}\nu_{k}\leq\inf_{k\in\mathbb{N}}\gamma(\overline{\theta_{X^{\mathrm{in}}}},k)\leq\gamma(\overline{\theta_{X^{\mathrm{in}}}},0)<+\infty$
as $t\mapsto\gamma(s,t)$ is decreasing for all $s\in\mathbb{R}$. This proves that $\nu_{\rm opt}<+\infty$.
Suppose that $\nu=(\nu_{k})_{k\in\mathbb{N}}$ has a finite supremum. From the first statement of Th. 1, $\mathbb{\Gamma}(\nu)$ is nonempty and let $(h,\beta)$ be in $\mathbb{\Gamma}(\nu)$. Then, we define for all $s\in\mathbb{R}$ and $t\in\mathbb{R}_{+}$,
$\gamma(s,t):=\min\{s,h(\beta^{t})\}$
and for all $x\in\mathbb{R}^{d}$,
$\theta(x):=\sup_{k\in\mathbb{N}}\varphi(T^{k}(x)).$
It is clear that $\gamma\in\mathcal{KL}_{\rm gen}$ as $\beta\in(0,1)$. Now by assumption, $\nu_{\rm opt}=\sup_{n\in\mathbb{N}}\nu_{n}\in\mathbb{R}$. Moreover, we have
$\nu_{\rm opt}=\sup_{n\in\mathbb{N}}\sup_{x\in X^{\mathrm{in}}}\varphi(T^{n}(x))=\sup_{x\in X^{\mathrm{in}}}\sup_{n\in\mathbb{N}}\varphi(T^{n}(x))=\overline{\theta_{X^{\mathrm{in}}}}.$
By definition of $\theta$, we have for all $x\in X^{\mathrm{in}}$, for all $k\in\mathbb{N}$, $\varphi(T^{k}(x))\leq\theta(x)$. By definition of $(h,\beta)$, we have for all $k\in\mathbb{N}$, $\nu_{k}\leq h(\beta^{k})$. This is the same as for all $x\in X^{\mathrm{in}}$ and for all $k\in X^{\mathrm{in}}$, $\varphi(T^{k}(x))\leq h(\beta^{k})$. We conclude that for all $x\in X^{\mathrm{in}}$ and for all $k\in\mathbb{N}$
$\varphi(T^{k}(x))\leq\min\{\theta(x),h(\beta^{k})\}.$
Note that for the function $\gamma$ constructed as above, for all $s\in\mathbb{R}$
$\inf_{t\geq 0}\gamma(s,t)=\inf_{t\geq 0}\min\{s,h(\beta^{t})\}=\min\{s,h(0)\}.$
This follows from the continuity of $h$, the fact $h$ is strictly increasing and $\beta\in(0,1)$.
@@
If moreover, ${\mathrm{R}}_{\nu}^{s}\neq\emptyset$, then from the second statement of Th 1, there exists $h\in\mathbb{\Gamma}_{\rm fun}(\nu)$ useful for $\nu$. We can define $\gamma$ using this function $h$. Now, let $k\in\mathcal{S}(\nu,h)$. As $\nu_{k}\leq\nu_{\rm opt}=\overline{\theta_{X^{\mathrm{in}}}}$, we conclude that
$\inf_{t\geq 0}\gamma(\overline{\theta_{X^{\mathrm{in}}}},t)=\min\{\overline{\theta_{X^{\mathrm{in}}}},h(0)\}=h(0)$
which ends the proof. ∎
@@
In the proof of Th. 2, we exploit Th. 1 which links the nonemptiness of the sets $\mathbb{\Gamma}(u)$, $\mathcal{S}(u,h)$ and ${\mathrm{R}}_{u}^{s}$. We can reformulate Th 2 as Corollary 1.
@@
In Th 2 and in Corollary 1, we construct a useful $\mathcal{KL}_{\rm gen}$-$\operatorname{Fin}\left(X^{\mathrm{in}}\right)$ upper bound for $(X^{\mathrm{in}},T,\varphi)$ from a useful $(h,\beta)\in\mathbb{\Gamma}(\nu)$. Now, we have to check whether it is possible to construct a useful $(h,\beta)\in\mathbb{\Gamma}(\nu)$ from a useful $\mathcal{KL}_{\rm gen}$-$\operatorname{Fin}\left(X^{\mathrm{in}}\right)$ upper bound for $(X^{\mathrm{in}},T,\varphi)$. Thus, we provide a positive response. The result is based on an extension of Sontag’s result for the class $\mathcal{KL}_{\rm gen}$. This permits to fix the value $\beta\in(0,1)$ to only consider problem on functions in $\Omega([0,1])$.
@@
Let $f:\mathbb{R}_{+}\to\mathbb{R}$ be a decreasing function and let $m\geq\inf_{\mathbb{R}_{+}}f$ ($m$ can be equal to -$\infty$ if $\inf_{\mathbb{R}_{+}}f$ is). There exists a function $g:\mathbb{R}_{+}\mapsto\mathbb{R}$ that is strictly decreasing and continuous such that:
@@
Now assume that $f(0)>m\geq\inf_{\mathbb{R}_{+}}f$. We denote for all couple of pairs $\{(x,y),(z,t)\}$ where $x,y,z,t\in\mathbb{R}$ by $\mathrm{Aff}\{(x,y),(z,t)\}$ the unique affine function $\ell:\mathbb{R}\mapsto\mathbb{R}$ such that $\ell(x)=y$ and $\ell(z)=t$. We define the functions $F_{0}$ and $F_{n}$ for all $n\in\mathbb{N}^{*}$ as follows:
@@
[…]
@@
The function $\sum_{n\in\mathbb{N}}F_{n}$ is clearly continuous and decreasing (as $f$ does) on $\mathbb{R}_{+}$. The function $\sum_{n\in\mathbb{N}}F_{n}$ is not strictly decreasing, but adding $x\mapsto(x+1)^{-1}$ makes $F$ strictly decreasing and continuous. We also have $f\leq\sum_{n\in\mathbb{N}}F_{n}$ because $f$ is decreasing. As $x\mapsto(x+1)^{-1}$ is strictly positive, we have $f\leq F$. Finally, we have $\inf_{\mathbb{R}_{+}}F=\inf_{\mathbb{R}_{+}}f$. Indeed, for all $n\in\mathbb{N}^{*}$ and all $x\in[n,n+1)$, $F_{n}\leq f(n-1)$, it implies that $\inf_{\mathbb{R}_{+}}F=\inf_{x\in\mathbb{R}_{+}}\inf_{n\in\mathbb{N}^{*}}F_{n}(x)\leq\inf_{n\in\mathbb{N}^{*}}f(n-1)=\inf_{\mathbb{R}_{+}}f$ (as $f$ decreases). The reverse inequality holds as $f\leq F$.
The function $F$ does not necessarily satisfy $\inf_{\mathbb{R}_{+}}F=m$. Hence, we define for all $x\in\mathbb{R}_{+}$ as
@@
Suppose that $m>\inf_{\mathbb{R}_{+}}f$. The function $g$ is clearly strictly decreasing and continuous and satisfies $f\leq g$ as $F$ does. As $m>\inf_{\mathbb{R}_{+}}f=\inf_{\mathbb{R}_{+}}F$, there exists $x_{0}\in\mathbb{R}_{+}$ such that $F(x)<m$ for all $x\geq x_{0}$ and then $g(x)=(x+1)^{-1}+m$ for all $x\geq x_{0}$. This implies that $\inf_{\mathbb{R}_{+}}g=m$. ∎
@@
By definition, $t\mapsto\gamma(s,t)$ is decreasing and applying Lemma 1, there exists a function $\sigma:\mathbb{R}_{+}\mapsto\mathbb{R}$ strictly decreasing and continuous such that $\gamma(s,t)\leq\sigma(t)$ for all $t\geq 0$ and $\inf_{\mathbb{R}_{+}}\sigma=m$. Let us define for all $r\in(0,1]$, $h(r)=\sigma(-\ln r)$ and $\lim_{x\to 0^{+}}h(x)=\lim_{y\to+\infty}\sigma(y)=m$. Then, defining $h(0)=m$ makes $h$ continuous at 0. Finally, $h$ belongs to $\Omega([0,1])$. ∎
Proof of Th. 3.
@@
By assumption, there exists a $\mathcal{KL}_{\rm gen}$-$\operatorname{Fin}\left(X^{\mathrm{in}}\right)$ upper bound $(\gamma,\theta)$ for $(X^{\mathrm{in}},T,\varphi)$. Let $x\in X^{\mathrm{in}}$ and $k\in\mathbb{N}$. This implies that $\varphi(T^{k}(x))\leq\gamma(\overline{\theta_{X^{\mathrm{in}}}},k)$. By Lemma 2, there exists $h\in\Omega([0,1])$ such that $\gamma(\overline{\theta_{X^{\mathrm{in}}}},k)\leq h(e^{-k})$ for all $k\in\mathbb{N}$. This leads to $(h,e^{-1})\in\mathbb{\Gamma}(\nu)$. If there exists $k\in\mathbb{N}$ such that $\inf_{t\in\mathbb{R}}\gamma(\overline{\theta_{X^{\mathrm{in}}}},t)<\nu_{k}$, we can choose, as it is stated in Lemma 2, $h$ such that $h(0)=m\in(\inf_{t\in\mathbb{R}}\gamma(\overline{\theta_{X^{\mathrm{in}}}},t),\nu_{k})$. ∎
Example 3 (An element of $\mathbb{\Gamma}(\omega)$ from $(\gamma_{\overline{\varepsilon}},\theta_{\ell})$).
From Th 3, we can build an element of $\mathbb{\Gamma}(\omega)$ from the $\mathcal{KL}_{\rm gen}$-$\operatorname{Fin}\left(X_{\ell}\right)$ upper bound obtained in Example 2. The situation is simpler in Example 2 as we get directly a scalar in $(0,1)$ from the expression of $\gamma_{\overline{\varepsilon}}$.
Let
$h:[0,1]\ni x\mapsto\left\{\begin{array}[]{lr}\displaystyle{\gamma_{\overline{\varepsilon}}\left(\overline{{\theta_{\ell}}_{X_{\ell}}},\dfrac{\ln(x)}{\ln(\beta)}\right)}&\text{ if }x\in(0,1]\\ \dfrac{\lambda_{2}}{c}&x=0\end{array}\right.\text{ and }\beta=\left|\dfrac{\lambda_{1}}{\lambda_{2}}\right|$
where $\gamma_{\varepsilon}$ is defined in (28), $\overline{\varepsilon}$ and $\theta_{\ell}$ are defined in Example 2. As $t$ tends to $\infty$, we have $\overline{{\theta_{\ell}}_{X_{\ell}}}<-|\lambda_{1}/\lambda_{2}|^{t}-\overline{\varepsilon}$ for all sufficiently large $t$. It follows that $\lim_{x\mapsto 0^{+}}h(x)=\lim_{z\mapsto 0^{+}}H(\overline{{\theta_{\ell}}_{X_{\ell}}},z)=\lambda_{2}/c$. Thus, the function $h$ is continuous at 0. From (31), we can deduce that $(h,\beta)\in\mathbb{\Gamma}(\omega)$. Furthermore, we have, $h([0,1])=[\lambda_{2}/c,\gamma_{\overline{\varepsilon}}(\overline{{\theta_{\ell}}_{X_{\ell}}},0)]$ that contains all terms $\omega_{k}$ greater than the limit superior of $\omega$ and
$\lambda_{2}/c,\gamma_{\overline{\varepsilon}}(\overline{{\theta_{\ell}}_{X_{\ell}}},0)]\ni y\mapsto\displaystyle{{h}_{\left([0,1]\right)}^{-1}}(y)=\left\{\begin{array}[]{lr}\dfrac{\overline{{\theta_{\ell}}_{X_{\ell}}}(\lambda_{2}-cy)}{cy-\lambda_{1}}&\text{ if }y\leq\dfrac{\overline{\varepsilon}\lambda_{1}-\overline{{\theta_{\ell}}_{X_{\ell}}}(\lambda_{2}-\lambda_{1})}{c\overline{\varepsilon}}\\ \dfrac{\overline{\varepsilon}(cy-c\overline{{\theta_{\ell}}_{X_{\ell}}}-c\overline{\varepsilon}-\lambda_{2})}{c\overline{\varepsilon}+\lambda_{2}-\lambda_{1}}&\text{ if }y>\dfrac{\overline{\varepsilon}\lambda_{1}-\overline{{\theta_{\ell}}_{X_{\ell}}}(\lambda_{2}-\lambda_{1})}{c\overline{\varepsilon}}\end{array}\right.$.
Then we can use Formula (4) that becomes, if $\omega_{k}>\lambda_{2}/c$:
$\dfrac{\ln(\displaystyle{{h}_{\left([0,1]\right)}^{-1}}(\omega_{k}))}{\ln\left(\left|\dfrac{\lambda_{1}}{\lambda_{2}}\right|\right)}.$
Applying those computations to the numerical data of Subsection 3.4, we get, as $\overline{{\theta_{\ell}}_{X_{\ell}}}=-11/19$
$\left[4,\dfrac{11891}{76}\right]\ni y\mapsto\displaystyle{{h}_{\left([0,1]\right)}^{-1}}(y)=\left\{\begin{array}[]{lr}-\dfrac{11}{19}\dfrac{(2-0.5y)}{0.5y+1}&\text{ if }y\leq 86\\ \\ \dfrac{y}{153}-\dfrac{263}{11628}&\text{ if }y>86\end{array}\right.$.
Recall that $\omega_{0}=\varphi_{\ell}(x^{ul})=40/5=8$ and then
$\dfrac{\ln(\displaystyle{{h}_{\left([0,1]\right)}^{-1}}(8))}{-\ln(2)}=\dfrac{\ln(22/57)}{-\ln(2)}\approx 1.3735$
This means that the maximum of terms $(\omega_{k})_{k}$ is either $\omega_{0}$ or $\omega_{1}$. Now as $\omega_{1}=\varphi_{\ell}(Lx^{lu})=42$. This implies that $\omega_{\rm opt}=42$. Moreover, we find that
$\dfrac{\ln(\displaystyle{{h}_{\left([0,1]\right)}^{-1}}(42))}{-\ln(2)}=\dfrac{\ln\left(\dfrac{11\times 19}{19\times 22}\right)}{-\ln(2)}=1.$
4.2 Summary of the $\mathcal{KL}_{\rm gen}$ extension